# You Sound a Little Tense: L2 Tailored Clear TTS Using Durational Vowel Properties

### v1

@@ 4 Discussion and future work
Further, through explorations of the results on the participant level in [22], we found that the duration mechanism was not clearly universal; Rather, it appears to be very strong in certain participants and weak or non-existent in others, which echoes work showing inter-individual differences in the weighting of L2 acoustic cues in speech perception [33]. This suggests avenues for future work, which could harness pronunciation data to further customize TTS to individual listeners. Moreover, since our work used force choice to limit participant responses, a topic of future work is understanding how increasing clarity in a target word affects the rest of the phrase and how robust this duration mechanism is to broader linguistic contexts. Additionally, since our participants had a relatively high level of English proficiency, future work should explore “L2 clarity TTS” with participants of lower proficiency levels, which could explore the modification of words that are less likely to be in the vocabulary of L2 speakers (e.g., “cooed” vs. “could”), using individual vocabularies as a factor predicting perception. Lastly, even more fine-grained control over a TTS [34] could aid in uncovering other possible mechanisms to aid L2 perception, specifically for how formant control and duration control can be combined.
@@ 5 Conclusions
This study resulted in multiple interesting findings. First, we provided a “clarity mode” as an open-source addition to Matcha-TTS and confirmed that, by applying a stretch to tense vowels for difficult target words, L2 speakers’ transcription errors for these words were reduced. Indeed, emphasizing all target words or simply slowing the whole phrase without consideration of the linguistic properties of the vowel reduced the transcription performance of L2 speakers (from French L1 backgrounds), who primarily use duration to determine the difference between English tense and lax vowels. Moreover, we confirm prior findings that slowing down an entire phrase can be seen as less respectful and encouraging to L2 listeners.
@@
Overall, these results have important consequences for TTS assessments, especially for L2 speakers. We cannot simply rely on the same methods used for L1 speaker TTS assessments. Neither user self-rated intelligibility assessments nor automatic systems reflect the true accuracy of difficult vowel perception in L2 speakers, even for those with high proficiency as in our sample, and improved accuracy does not necessarily reflect positive perceptions of the voice (as for “stretch”). Researchers must use both objective and subjective assessments with human participants to ensure they are building inclusive and accessible speech synthesis systems.

### v2

@@ 4 Discussion and future work
Further, through explorations of the results on the participant level in [22], we found that the duration mechanism was not clearly universal; Rather, it appears to be very strong in certain participants and weak or non-existent in others, which echoes work showing inter-individual differences in the weighting of L2 acoustic cues in speech perception [33]. This suggests avenues for future work, which could harness pronunciation data to further customize TTS to individual listeners. Moreover, since our work used force choice to limit participant responses, a topic of future work is understanding how increasing clarity in a target word affects the rest of the phrase and how robust this duration mechanism is to broader linguistic contexts. Additionally, since our participants had a relatively high level of English proficiency, future work should explore ``L2 clarity TTS'' with participants of lower proficiency levels, which could explore the modification of words that are less likely to be in the vocabulary of L2 speakers (e.g., ``cooed'' vs. ``could''), using individual vocabularies as a factor predicting perception. Lastly, even more fine-grained control over a TTS [34] could aid in uncovering other possible mechanisms to aid L2 perception, specifically for how formant control and duration control can be combined.
@@ 5 Conclusions
This study resulted in multiple interesting findings. First, we provided a ``clarity mode'' as an open-source addition to Matcha-TTS and confirmed that, by applying a stretch to tense vowels for difficult target words, L2 speakers' transcription errors for these words were reduced. Indeed, emphasizing all target words or simply slowing the whole phrase without consideration of the linguistic properties of the vowel reduced the transcription performance of L2 speakers (from French L1 backgrounds), who primarily use duration to determine the difference between English tense and lax vowels. Moreover, we confirm prior findings that slowing down an entire phrase can be seen as less respectful and encouraging to L2 listeners.
@@
Overall, these results have important consequences for TTS assessments, especially for L2 speakers. We cannot simply rely on the same methods used for L1 speaker TTS assessments. Neither user self-rated intelligibility assessments nor automatic systems reflect the true accuracy of difficult vowel perception in L2 speakers, even for those with high proficiency as in our sample, and improved accuracy does not necessarily reflect positive perceptions of the voice (as for ``stretch''). Researchers must use both objective and subjective assessments with human participants to ensure they are building inclusive and accessible speech synthesis systems.